# Large Language Model-Brained GUI Agents: A Survey

### v9

@@ 6 LLM-Brained GUI Agent Framework
TABLE VII: Overview of LLM-brained GUI agent frameworks on web platforms (Part I).
@@
TABLE VIII: Overview of LLM-brained GUI agent frameworks on web platforms (Part II).
@@ 6.1 Web GUI Agents
Advancements in web GUI agents have led to significant strides in automating complex tasks within diverse and dynamic web environments. Recent frameworks have introduced innovative approaches that leverage multimodal inputs, predictive modeling, and task-specific optimizations to enhance performance, adaptability, and efficiency. In this subsection, we first summarize key web GUI agent frameworks in Tables VII and VIII, then delve into representative frameworks, highlighting their unique contributions and how they collectively push the boundaries of web-based GUI automation.
@@
TABLE IX: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part I).
@@
TABLE X: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part II).
@@ 6.2 Mobile GUI Agents
The evolution of mobile GUI agents has been marked by significant advancements, leveraging multimodal models, complex architectures, and adaptive planning to address the unique challenges of mobile environments. These agents have progressed from basic interaction capabilities to sophisticated systems capable of dynamic, context-aware operations across diverse mobile applications. We first provide an overview of mobile GUI agent frameworks in Tables IX and X.
@@
TABLE XI: Overview of LLM-brained GUI agent frameworks on computer platforms.
@@ 6.3 Computer GUI Agents
Computer GUI agents have evolved to offer complex automation capabilities across diverse operating systems, addressing challenges such as cross-application interaction, task generalization, and high-level task planning. They have led to the development of sophisticated frameworks capable of handling complex tasks across desktop environments. These agents have evolved from simple automation tools to intelligent systems that leverage multimodal inputs, advanced architectures, and adaptive learning to perform multi-application tasks with high efficiency and adaptability. We provide an overview of computer GUI agent frameworks in Table XI.
@@
TABLE XII: Overview of LLM-brained cross-platform GUI agent frameworks.
@@ 6.4 Cross-Platform GUI Agents
Cross-platform GUI agents have emerged as versatile solutions capable of interacting with various environments, from desktop and mobile platforms to more complex systems. These frameworks prioritize adaptability and efficiency, leveraging both lightweight models and multi-agent architectures to enhance cross-platform operability. In this subsection, we first We overview cross-platform GUI agent frameworks in Table XII, then explore key frameworks that exemplify the advancements in cross-platform GUI automation.
@@ 7.1.2 Collection Pipeline
TABLE XIII: Overview of datasets for optimizing LLMs tailored for web GUI agents.
@@ 7.2 Web Agent Data
Web-based GUI agents demand datasets that capture the intricate complexity and diversity of real-world web interactions. These datasets often encompass varied website structures, including DOM trees and HTML content, as well as multi-step task annotations that reflect realistic user navigation and interaction patterns. Developing agents that can generalize across different websites and perform complex tasks requires comprehensive datasets that provide rich contextual information. We provide an overview of web-based GUI agents dataset in Table XIII.
@@
TABLE XIV: Overview of datasets for optimizing LLMs tailored for mobile GUI agents (Part I).
@@
TABLE XV: Overview of datasets for optimizing LLMs tailored for mobile GUI agents (Part II).
@@ 7.3 Mobile Agent Data
Mobile platforms are critical for GUI agents due to the diverse range of apps and unique user interactions they involve. To develop agents that can effectively navigate and interact with mobile interfaces, datasets must offer a mix of single and multi-step tasks, focusing on natural language instructions, UI layouts, and user interactions. We first overview mobile GUI agents dataset in Tables XIV and XV.
@@
TABLE XVI: Overview of datasets for optimizing LLMs tailored for computer GUI agents.
@@ 7.4 Computer Agent Data
In contrast to mobile and web platforms, the desktop domain for GUI agents has relatively fewer dedicated datasets, despite its critical importance for applications like productivity tools and enterprise software. However, notable efforts have been made to support the development and evaluation of agents designed for complex, multi-step desktop tasks. We show related dataset for computer GUI agents in Table XVI.
@@
TABLE XVII: Overview of datasets for optimizing LLMs tailored for cross-platform GUI agents.
@@ 7.5 Cross-Platform Agent Data
Cross-platform datasets play a pivotal role in developing versatile GUI agents that can operate seamlessly across mobile, computer, and web environments. Such datasets support generalizability and adaptability, enabling agents to handle varied interfaces and tasks in real-world applications. We provide an overview of related dataset for cross-platform GUI agents in Table XVII.

### v11

@@ 6 LLM-Brained GUI Agent Framework
TABLE X: Overview of LLM-brained GUI agent frameworks on web platforms (Part I).
@@
TABLE XI: Overview of LLM-brained GUI agent frameworks on web platforms (Part II).
@@
TABLE XII: Overview of LLM-brained GUI agent frameworks on web platforms (Part III).
Agent Platform Perception Action Model Architecture Highlight Link
@@
AgentSymbiotic [289] Web Accessible tree structure of web elements Standard UI operations Large LLMs: GPT-4o, Claude-3.5. Small LLMs: LLaMA-3 [90], DeepSeek-R1 [495] Multi-agent iterative architecture Introduces an iterative, symbiotic learning process between large and small LLMs for web automation. Enhances both data synthesis and task performance through speculative data synthesis, multi-task learning, and privacy-preserving hybrid modes. /
LiteWebAgent [292] Web DOM, Screenshots Standard UI operations, Playwright script Any LLM and MLLM Single-agent First open-source, production-ready web agent integrating tree search for multi-step task execution. https://github.com/PathOnAI/LiteWebAgent
ECLAIR [291] Web Screenshots Standard UI operations GPT-4V, GPT-4o, CogAgent [496] Single-agent architecture Eliminates the high setup costs, brittle execution, and burdensome maintenance associated with traditional RPA by learning from video and text documentation. https://github.com/HazyResearch/eclair-agents
Dammu et al., [293] Web DOM elements, Webpage accessibility attributes Standard UI operations Not specified Single-agent architecture User-aligned task execution, where the agent adapts to individual user preferences in an ethical manner. /
Plan-and-Act [294] Web HTML Standard UI operations LLaMA-3.3-70B-Instruct [90] Two-stage modular architecture: PLANNER + EXECUTOR Decouples planning from execution in LLM-based GUI agents and introduces a scalable synthetic data generation pipeline to fine-tune each component /
SkillWeaver [295] Web GUI screenshots and Accessibility Tree Standard UI operations and high-level skill APIs GPT-4o Single-agent Introduces a self-improvement framework for web agents that autonomously discover, synthesize, and refine reusable skill APIs through exploration https://github.com/OSU-NLP-Group/SkillWeaver
ASI [296] Web Webpage Accessibility Tree Standard GUI actions Claude-3.5-Sonnet Single-agent Introduces programmatic skills that are verified through execution to ensure quality and are used as callable actions to improve efficiency https://github.com/zorazrw/agent-skill-induction
Rollback Agent [297] Web Accessibility trees Standard GUI actions Multi-agent architecture Multi-module, ReAct-inspired agent architecture Introduces a modular rollback mechanism that enables multi-step rollback to avoid dead-end states /
@@ 6.1 Web GUI Agents
Figure 22: Comparison of RPA and agent based automation. Figure adapted from [291].
Advancements in web GUI agents have led to significant strides in automating complex tasks within diverse and dynamic web environments. Recent frameworks have introduced innovative approaches that leverage multimodal inputs, predictive modeling, and task-specific optimizations to enhance performance, adaptability, and efficiency. In this subsection, we first summarize key web GUI agent frameworks in Tables X, XI and XII, then delve into representative frameworks, highlighting their unique contributions and how they collectively push the boundaries of web-based GUI automation.
@@
ECLAIR [291] represents a pioneering application that replaces traditional RPA with a foundation model-powered GUI agent for enterprise automation. Unlike conventional RPA, which relies on manually programmed rules and rigid scripts, ECLAIR dynamically learns workflows from video demonstrations and textual SOPs (Standard Operating Procedures), significantly reducing setup time and improving adaptability. It operates on enterprise web applications, leveraging GPT-4V and CogAgent [496] to perceive GUI elements, plan actions, and execute workflows, and validate automatically. By eliminating the high maintenance costs and execution brittleness of RPA, ECLAIR introduces a more flexible and scalable approach to GUI automation. We show a comparison of such agent-based vs. RPA automation in Figure 22. This work establishes an important foundation for LLM-powered GUI automation, demonstrating how multimodal foundation models can bridge the gap between process mining, RPA, and fully autonomous enterprise workflows.
@@
TABLE XIII: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part I).
@@
TABLE XIV: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part II).
@@
TABLE XV: Overview of LLM-brained GUI agent frameworks on mobile platforms (Part III).
Agent Platform Perception Action Model Architecture Highlight Link
@@
Mobile-Agent-V [316] Mobile Android Video guidance, XML hierarchy Standard UI operations GPT-4o Multi-agent system Introduces video-guided learning, allowing the agent to acquire operational knowledge efficiently. https://github.com/X-PLUG/MobileAgent
MobileSteward [317] Mobile Android XML layouts, Screenshots Standard UI interactions, Code execution GPT-4V, GPT-4o App-oriented multi-agent framework Introduces an app-oriented multi-agent framework with self-evolution, overcoming the complexity of cross-app interactions by dynamically recruiting specialized agents. https://github.com/XiaoMi/MobileSteward
AppAgentX [258] Mobile Android Screenshots Standard UI operations GPT-4o Single-agent architecture Introduces an evolutionary mechanism that enables dynamic learning from past interactions and replaces inefficient low-level operations with high-level actions. https://appagentx.github.io/
CHOP [318] Mobile Android Screenshots Standard UI operations GPT-4o Multi-agent architecture Introduces a basis subtask framework, where subtasks are predefined based on human task decomposition patterns, ensuring better executability and efficiency. https://github.com/Yuqi-Zhou/CHOP
OS-Kairos [319] Mobile Android GUI screenshots Standard UI operations OS-Atlas-Pro-7B and GPT-4o Single-agent with critic-in-the-loop design Introduces an adaptive interaction framework where each GUI action is paired with a confidence score, dynamically deciding between autonomous execution and human intervention https://github.com/Wuzheng02/OS-Kairos
V-Droid [320] Mobile Android Android Accessibility Tree Standard UI operations LLaMA-3.1-8B-Instruct [90] Verifier-Driven Single-Agent Architecture Introduces a novel verifier-driven architecture where the LLM does not generate actions directly but instead scores and selects from a finite set of extracted actions, improving task success rates and significantly reducing latency /
LearnAct [321] Mobile Android GUI screenshots, UI trees, and demonstration trajectories Standard GUI actions Gemini-1.5-Pro, UI-TARS-7B-SFT, Qwen2-VL-7B Multi-agent Introduces a structured, demonstration-based learning pipeline for mobile GUI agents. It addresses long-tail generalization via few-shot demonstrations, achieving substantial performance gains on complex real-world mobile tasks https://lgy0404.github.io/LearnAct
@@ 6.2 Mobile GUI Agents
The evolution of mobile GUI agents has been marked by significant advancements, leveraging multimodal models, complex architectures, and adaptive planning to address the unique challenges of mobile environments. These agents have progressed from basic interaction capabilities to sophisticated systems capable of dynamic, context-aware operations across diverse mobile applications. We first provide an overview of mobile GUI agent frameworks in Tables XIII XIV and XV.
@@
TABLE XVI: Overview of LLM-brained GUI agent frameworks on computer platforms (Part I)..
@@
UFO2 [330] Windows desktops GUI screenshots and textual control properties list Unified GUI–API action layer GPT-4o (and GPT-4V, o1, Gemini-Flash); Vision grounding via OmniParser-v2 Centralized HostAgent with application-specialized AppAgents Transforms a conventional CUA into an OS-native, pluggable AgentOS with deep Windows integration, hybrid GUI–API actions, vision + UIA perception, speculative multi-action planning, retrieval-augmented knowledge, and a non-intrusive PiP virtual desktop https://github.com/microsoft/UFO/
@@
TABLE XVII: Overview of LLM-brained GUI agent frameworks on computer platforms (Part II).
Agent Platform Perception Action Model Architecture Highlight Link
@@
PC-Agent [326] Windows computers UI tree, Screenshots Standard UI operations GPT-4o Hierarchical Multi-Agent PC-Agent’s hierarchical multi-agent design enables efficient decomposition of complex PC tasks. Its Active Perception Module enhances fine-grained GUI understanding by combining accessibility structures, OCR, and intention grounding. https://github.com/X-PLUG/MobileAgent/tree/main/PC-Agent
PwP [327] VSCode-based IDE in Computers Screenshots, File system access, Terminal outputs Standard UI interactions, File operations, Bash commands, Tools in VSCode GPT-4o, Claude-3.5 Sonnet, Gemini-1.5 Single-agent architecture Shifts software engineering agents from API-based tool interactions to direct GUI-based computer use, allowing agents to interact with an IDE as a human developer would. https://programmingwithpixels.com
COLA [328] Windows computers GUI structure, properties and screenshots Standard UI operations and system APIs GPT-4o Hierarchical Multi-Agent A dynamic task scheduling mechanism with a plug-and-play agent pool, enabling adaptive handling of GUI tasks https://github.com/Alokia/COLA-demo
STEVE [329] Windows Desktop GUI screenshots and A11y Tree Standard UI operations Qwen2-VL [232] and GPT-4o Single-agent Introduces a scalable step verification pipeline using GPT-4o to generate binary labels for agent actions, and applies KTO optimization to incorporate both positive and negative actions into agent learning https://github.com/FanbinLu/STEVE
@@ 6.3 Computer GUI Agents
Computer GUI agents have evolved to offer complex automation capabilities across diverse operating systems, addressing challenges such as cross-application interaction, task generalization, and high-level task planning. They have led to the development of sophisticated frameworks capable of handling complex tasks across desktop environments. These agents have evolved from simple automation tools to intelligent systems that leverage multimodal inputs, advanced architectures, and adaptive learning to perform multi-application tasks with high efficiency and adaptability. We provide an overview of computer GUI agent frameworks in Table XVI and XVII.
@@
Figure 24: The comparison of traditional CUAs and the Desktop AgentOS UFO2. Figure adapted from the original paper.
UFO2 [330], the successor to UFO, elevates GUI automation from a vision-only prototype to a deeply integrated, Windows-native AgentOS (Figure 24). It coordinates tasks through a centralized HostAgent, which delegates subtasks to application-specialized AppAgents. A hybrid perception pipeline that fuses Windows UI Automation (UIA) metadata with OmniParser-v2 visual grounding delivers robust control identification even for custom widgets. Via a unified GUI–API action layer, AppAgents preferentially invoke high-level application APIs and fall back to pixel-level clicks only when necessary, cutting both latency and brittleness. A picture-in-picture virtual desktop cleanly isolates agent execution from the user’s main session, enabling non-intrusive multitasking. Runtime performance is further boosted by retrieval-augmented help documents and execution logs, coupled with speculative multi-action planning that executes several steps per single LLM invocation. Tested on 20+ real Windows applications, UFO2 exceeds Operator [504] and other CUAs by more than 10 percentage points in success rate while halving LLM calls. Because the framework is model-agnostic, swapping GPT-4o for a stronger LLM such as o1 yields additional gains without code changes.
@@
In the emerging field of LLM-powered GUI agents for desktop environments, Programming with Pixels (PwP) [327] introduces a compelling alternative to traditional tool-based software engineering agents. Rather than relying on predefined API calls, PwP enables agents to interact directly with an IDE using visual perception, keyboard inputs, and mouse clicks, mimicking the way human developers operate within an IDE. This approach allows for generalization beyond predefined APIs, providing a highly expressive environment where agents can execute a wide range of software engineering tasks, including debugging, UI generation, and code editing. Evaluations conducted on PwP-Bench demonstrate that computer-use agents, despite lacking direct access to structured APIs, can match or even surpass traditional tool-based approaches in certain scenarios.
@@
TABLE XVIII: Overview of LLM-brained cross-platform GUI agent frameworks.
@@
CollabUIAgents [334] Mobile Android, Web Screenshots, UI trees Standard UI operations Qwen2-7B [232], GPT-4 Multi-agent system A multi-agent reinforcement learning framework that introduces a Credit Re-Assignment (CR) strategy, using LLMs instead of environment-specific rewards to enhance performance and generalization. https://github.com/THUNLP-MT/CollabUIAgents
Agent S2 [339] Ubuntu, Windows, Android GUI screenshot Standard UI operations and system APIs Claude-3.7-Sonnet, Claude-3.5-Sonnet, GPT-4o (for Manager and Worker roles), UI-TARS-72B-DPO, Tesseract OCR, and UNO (for grounding experts) Compositional multi-agent architecture with a Manager for planning, a Worker for execution, and a Mixture of Grounding experts Features a Mixture of Grounding technique and Proactive Hierarchical Planning, enabling more accurate grounding and adaptive replanning in long-horizon tasks https://github.com/simular-ai/Agent-S
GuidNav [340] Android and Web GUI screenshots Standard UI operations and system APIs GPT-4o, Gemini 2.0 Flash, Qwen-VL-Plus Single-agent Introduces a novel process reward model that provides fine-grained, step-level feedback to enhance GUI task accuracy and success /
@@ 6.4 Cross-Platform GUI Agents
Cross-platform GUI agents have emerged as versatile solutions capable of interacting with various environments, from desktop and mobile platforms to more complex systems. These frameworks prioritize adaptability and efficiency, leveraging both lightweight models and multi-agent architectures to enhance cross-platform operability. In this subsection, we first We overview cross-platform GUI agent frameworks in Table XVIII, then explore key frameworks that exemplify the advancements in cross-platform GUI automation.
@@
Agent S2 [339] builds upon its predecessor, Agent S [322], by introducing a hierarchical and compositional framework for GUI agents that integrates generalist models with specialized grounding modules. Departing from monolithic architectures, it employs a Mixture of Grounding (MoG) strategy to delegate fine-grained grounding tasks to expert modules, and adopts Proactive Hierarchical Planning (PHP) to dynamically revise action plans based on evolving observations. Relying solely on GUI screenshots, Agent S2 generalizes effectively across Ubuntu, Windows, and Android platforms. It demonstrates strong scalability and consistently outperforms larger monolithic models by strategically distributing cognitive responsibilities. The design of Agent S2 underscores the advantages of modular architectures for handling long-horizon, high-fidelity GUI interactions.
@@ 7.1.2 Collection Pipeline
TABLE XIX: Overview of datasets for optimizing LLMs tailored for web GUI agents.
@@
Explorer [290] Web Popular URLs with systematic web exploration by LLMs Textual task descriptions, Action sequences, GUI screenshots, Accessibility trees, HTML content 94K successful web trajectories, 49K unique URLs, 720K screenshots Generated by a multi-agent LLM pipeline Largest-scale web trajectory dataset to date; dynamically explores web pages to create contextually relevant tasks /
InSTA [345] Web Automatically generated by LLMs across 1M websites from Common Crawl Web navigation tasks in natural language, task plans and action sequences, HTML-based observations converted to markdown, and evaluations from LLM-based judges 150,000 tasks across 150,000 websites Generated by LLMs using the Playwright API and filtered by LLM-based judges Presents a fully automated three-stage data generation pipeline—task generation, action execution, and evaluation—using only language models without any human annotations https://data-for-agents.github.io
@@ 7.2 Web Agent Data
Web-based GUI agents demand datasets that capture the intricate complexity and diversity of real-world web interactions. These datasets often encompass varied website structures, including DOM trees and HTML content, as well as multi-step task annotations that reflect realistic user navigation and interaction patterns. Developing agents that can generalize across different websites and perform complex tasks requires comprehensive datasets that provide rich contextual information. We provide an overview of web-based GUI agents dataset in Table XIX.
@@
InSTA [345] is an Internet-scale dataset for training GUI-based web agents, generated entirely through an automated LLM pipeline without human annotations. It covers 150k diverse websites sourced from Common Crawl and includes rich web navigation tasks, trajectories in Playwright API calls, and evaluations using LLM-based judges. The dataset highlights strong generalization capabilities and data efficiency, significantly outperforming human-collected datasets like Mind2Web [213] and WebLINX [342] in zero-shot and low-resource settings. InSTA represents a key advancement in scalable data curation for LLM-powered GUI agents, offering unprecedented coverage across real-world web interfaces.
@@
TABLE $X$: Overview of datasets for optimizing LLMs tailored for mobile GUI agents (Part I).
@@
GUI-Xplore [359] Mobile Android Combination of automated exploration and manual design Exploration videos, textual tasks, QA pairs, view hierarchies, GUI screenshots, action sequences, and GUI transition graphs 312 apps, 115 hours of video, 32,569 QA pairs, 41,293 actions, about 200 pages per app Automated and human exploration Introduces an exploration-based pretraining paradigm that provides rich app-specific priors through video data https://github.com/921112343/GUI-Xplore
TABLE XXI: Overview of datasets for optimizing LLMs tailored for mobile GUI agents (Part II).
@@
FedMABench [445] Android Mobile AndroidControl [506], AITW [352] Textual task descriptions, action sequences, and GUI screenshots 6 dataset series with over 30 subsets Inferred from existing Android datasets The first dataset designed to benchmark federated mobile GUI agents https://github.com/wwh0411/FedMABench
@@ 7.3 Mobile Agent Data
Mobile platforms are critical for GUI agents due to the diverse range of apps and unique user interactions they involve. To develop agents that can effectively navigate and interact with mobile interfaces, datasets must offer a mix of single and multi-step tasks, focusing on natural language instructions, UI layouts, and user interactions. We first overview mobile GUI agents dataset in Tables XX and XXI.
@@
TABLE XXII: Overview of datasets for optimizing LLMs tailored for computer GUI agents.
@@
DeskVision [362] Windows, macOS, and Linux desktops Internet GUI screenshots with annotated bounding boxes for UI elements and detailed region captions 54,855 screenshots with 303,622 UI element annotations UI elements detected using OmniParser and PaddleOCR The first large-scale, open-source dataset focusing on real-world desktop GUI scenarios across operating systems /
@@ 7.4 Computer Agent Data
In contrast to mobile and web platforms, the desktop domain for GUI agents has relatively fewer dedicated datasets, despite its critical importance for applications like productivity tools and enterprise software. However, notable efforts have been made to support the development and evaluation of agents designed for complex, multi-step desktop tasks. We show related dataset for computer GUI agents in Table XXII.
@@
TABLE XXIII: Overview of datasets for optimizing LLMs tailored for cross-platform GUI agents (Part I).
@@
TABLE XXIV: Overview of datasets for optimizing LLMs tailored for cross-platform GUI agents (Part II).
@@
Navi-plus [371] Web and Android AndroidControl [506] and Mind2Web [213] Task descriptions, GUI action trajectories, low-level step instructions, screenshots, and follow-up ASK/SAY interaction pairs / LLM-automated with human validation Introduces a Self-Correction GUI Navigation task featuring the novel ASK action for recovering missing information /
Explorer [370] Web and Android Automated traversal of real websites and Android apps UI screenshots, bounding boxes of interactable elements, screen similarity labels, and user actions KhanAcademy (Web): 2,841 interactables, 378 screen similarity samples; Spotify (Android): 1,207 interactables, 451 screen similarity samples Automated tools, HTML parsing, Accessibility Tree Platform-independent, supports auto-labeling, and enables trace recording and voice-controlled GUI navigation https://github.com/varnelis/Explorer
@@ 7.5 Cross-Platform Agent Data
Cross-platform datasets play a pivotal role in developing versatile GUI agents that can operate seamlessly across mobile, computer, and web environments. Such datasets support generalizability and adaptability, enabling agents to handle varied interfaces and tasks in real-world applications. We provide an overview of related dataset for cross-platform GUI agents in Table XXIII and XXIV.